Evaluate the integral of e^(2lnx) dx


1
Expert's answer
2021-03-31T11:55:00-0400

"\int {{e^{2\ln x}}dx} = \int {{e^{\ln {x^2}}}dx} = \int {{x^2}} dx = \frac{{{x^3}}}{3} + C"

Answer: "\frac{{{x^3}}}{3} + C"